# 1]organization=Scuola Normale Superiore and INFN, addressline=Piazza dei Cavalieri 7, city=Pisa, postcode=56126, state=PI, country=Italy

### v2

@@ Abstract
After inflation.—After the inflation ends, the universe becomes almost flat. The interactions $\dot{\pi}\mathcal{O}_{\psi}$ and $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\psi}$ will freeze, since the energy scale becomes too small compared to the EFT of inflation scale $M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}$. However, even though the linear mixing interaction $\dot{\pi}\mathcal{O}_{\psi}$ remains small before it gets frozen, its nature as a relevant operator will drag our model into the gapped phase, creating bound states. They are natural dark matter candidates. We expect them to be standard SIMP-like dark matter. When we take the ratio between the dark matter mass and the freeze-out temperature to be the standard choice $x_{f}\equiv m_{\text{DM}}/T_{f}\sim 20$, we can believe that the lightest state dominates. By taking the freeze-out temperature to be below the QCD phase transition scale, the dark sector will remain in thermal equilibrium with the SM particles before freeze-out. The masses of the lowest lying states of pseudoscalars (PSs) and vector mesons (Vs) can be read from [13, 21]. The mass of the lightest vector mesons is 2.2 times the mass $m_{\text{PS}}$ of the lightest pseudoscalars, which means that the vector mesons are not the light vector mesons discussed in [22]. The pseudoscalars will dominate the relic abundance, since the vector mesons freeze out first. We can infer the mass scale of the dark pseudoscalars from the observed relic abundance. Since the bound states come from confinement, induced by a small relevant deformation, the information of interactions in the conformal phase has negligible impacts on the properties of bound states. It is worth mentioning that if the fermions have light masses in the UV [23, 24], the DM could become the dilaton forbidden dark matter, discussed in [25].
Based on the standard SIMP discussions in [26, 27], the dominant annihilation cross section is the $3\rightarrow 2$ self-annihilation cross section $\langle\sigma v^{2}\rangle_{3\rightarrow 2}$. It can be parametrized by an effective coupling $\alpha_{\text{eff}}$ in the way
$\langle\sigma v^{2}\rangle_{3\rightarrow 2}\equiv\alpha_{\text{eff}}^{3}/m_{\text{PS}}^{5}.$ (13)
Taking the DM to be non-relativistic at the freeze-out, the effective number of degrees of freedom is $g_{*}=10.75$.
[28] gives the expression for computing the relic abundance when the $3\rightarrow 2$ annihilation dominates, as
$\Omega_{\text{DM}}h^{2}=\frac{1.1\times 10^{-10}\text{ GeV}^{-2}}{\left(g_{*}^{3/2}\frac{m_{\text{PS}}^{2}}{M_{\text{pl}}}\int_{x_{f}}^{\infty}\frac{\text{d}x}{x^{5}}\left\langle\sigma v^{2}\right\rangle_{3\rightarrow 2}\right)^{1/2}}.$ (14)
To have the present-day observed value $\Omega_{\text{DM}}h^{2}=0.12$ [19], we found
$m_{\text{PS}}\sim\alpha_{\text{eff}}\times 30\text{ MeV}.$ (15)
If we take $\alpha_{\text{eff}}=1$ as in [26], $m_{\text{PS}}$ becomes $\sim 30\text{ MeV}$, which is a proper mass scale for SIMP dark matter.
@@
Discussion.—A simple gauge theory at the lower edge of the conformal window in the dark sector opens up new possibilities for beyond-Standard-Model (BSM) physics and dark matter candidates. The Dark Walker model has the potential to produce both the primordial non-Gaussianity, which is large enough to have the possibility of being detectable, and the observed relic abundance of dark matter. As expected, several parameters must still be fixed by hand, though their values remain within natural ranges. The non-Gaussianity signal is sensitive to the inflationary Hubble scale, the initial gauge coupling—which determines the starting energy scale of the merged fixed point—and the precise anomalous dimensions of the primary scalar operators in the conformal phase. These parameters may be constrained more accurately in the future through improved observations of inflationary parameters and lattice QCD studies of the spectrum in the $\text{SU}(3)$ theory with eight flavors. With such inputs, more precise predictions of the primordial non-Gaussianity and relic abundance can be achieved. This framework offers a concrete testing ground for the role of strongly coupled sectors in primordial cosmology.
@@ Acknowledgements
I express my thanks to Riccardo Barbieri, who encouraged me very helpfully while this work was in progress. Thanks to Raffaele Tito D’Agnolo, Janet Hung, Guilherme Leite Pimentel, and Alessandro Strumia for useful discussions and comments.

### v3

@@ Abstract
After inflation.—This section connects the inflationary phase to the present-day dark matter abundance. After the inflation ends, the universe becomes almost flat. The interactions $\dot{\pi}\mathcal{O}_{\psi}$ and $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\psi}$ will freeze, since the energy scale becomes too small compared to the EFT of inflation scale $M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}$. However, even though the linear mixing interaction $\dot{\pi}\mathcal{O}_{\psi}$ remains small before it freezes, its nature as a relevant operator will drag our model into the gapped phase. The gapped states are natural DM candidates. The masses of the lowest lying states of pseudoscalars (PSs) and vector mesons (Vs) can be read from the lattice QCD results [13, 23], where the mass of the lightest vector mesons $m_{\text{V}}$ is 2.2 times the mass of the lightest pseudoscalars $m_{\text{PS}}$. We will assume that the DM is made of the lightest pseudoscalars. We will use the freeze-in process to produce DM, mainly following the framework discussed in [24, 25]. It is worth mentioning that if the fermions have light masses in the UV [26, 27], the DM could become the dilaton forbidden dark matter, as discussed in [28].
If we assume that most of the energy from the inflaton is transferred to the SM after inflation, the SM is thermal at the reheating temperature $T_{R}$. Therefore, energy can be extracted from it and injected into the DS via a freeze-in process, due to the coupling between the DS operators and Higgs. [^1] We thank the anonymous reviewer for bringing this up. The model can be described by
$\mathcal{L}_{\text{SM}}+\mathcal{L}_{\text{CFT}}+\sum\lambda\frac{\mathcal{O}_{\text{H}}\mathcal{O}_{\Delta}}{\Lambda^{d}},\ d=d_{\text{SM}}+\Delta-4.$ (11)
Here $\mathcal{O}_{\text{SM}}$ is an SM singlet operator with dimension $d_{\text{SM}}$ and we will work in the regime of small dimensionless Wilson coefficient $\lambda\ll 1$. In the Higgs portal, the choice of scalar operators $\mathcal{O}_{\text{SM}}$ in the UV up to dimension four can be $|\text{H}|^{2}$ or $|D_{\mu}\text{H}|^{2}$. From [25] and also later, when $d>1/2$, the freeze-in processes are UV-dominated, and when $d<1/2$, the processes are IR-dominated. We assume the reheating temperature $T_{R}$ to be the highest temperature of the thermal bath, and it is much higher than the electroweak (EW) symmetry breaking scale $\sim 100\text{ GeV}$. The UV-dominating interaction will be $|D_{\mu}\text{H}|^{2}\mathcal{O}_{\psi}$, whose $d=1.8$. A natural choice for the effective coupling is $\Lambda^{\Delta}\equiv\Lambda_{\text{cw}}^{\Delta}$. The interaction $|D_{\mu}\text{H}|^{2}\mathcal{O}_{F}$, and the interactions between currents and stress tensors, e.g. $J_{\text{SM}}J_{\text{DS}}$, $T_{\text{SM}}T_{\text{DS}}$, etc., will be sub-dominant, because of the large scaling dimensions. As for the interaction $|\text{H}|^{2}\mathcal{O}_{F}$, the effective coupling is natural to be $\Lambda^{\Delta-2}\equiv\Lambda_{\text{cw}}^{\Delta}/\langle v\rangle^{2}$, where $\langle v\rangle$ is the vacuum expectation value (VEV) of the Higgs. This coupling is much suppressed in the UV regime. Nonetheless, the indirect interaction between the Higgs and $\mathcal{O}_{F}$ will also be sub-dominate and it is natural not to include this interaction in the UV freeze-in processes.
In the IR-dominant regime, the effective interaction will become
$\mathcal{L}_{\text{int}}=\lambda\frac{m^{2}}{\Lambda_{\text{cw}}^{\Delta}}\frac{v}{\sqrt{2}}h\mathcal{O}_{\text{CFT}},$ (12)
where $m$ is a mass parameter, which we will further assume to be the Higgs mass $m_{h}\sim 125\text{ GeV}$. The VEV of Higgs after EW symmetry breaking is $v\sim 246\text{ GeV}$. We assume that, in our scenario, the UV- and IR-dominant regimes are well-separated, and the energy density in the DS will remain far below the equilibrium with the SM. Thus the decay $\text{CFT}\rightarrow\text{SM}$ can be neglected. For now we will take $T_{R}/m_{h}>O(10)$, and we will see in the following that $T_{R}$ is actually lower-bounded by the permitted DM mass range.
We assume that the Dark Walker completely breaks the conformal symmetries and develops a mass gap $m_{\text{PS}}$ at scale $T_{D}\sim m_{\text{PS}}$. The SM can be treated as a free CFT with $m_{\text{PS}}\lesssim 100\text{ GeV}$ [24]. Thus the UV DS production can be treated as the Dark Walker CFT states from the SM CFT. From the classification of mass generation mechanisms in [14], we know that in the SMG case, the system is anomaly-free, and potentially the effective interactions between the pseudoscalars will be strongly coupled, since there is no small scale to suppress the interactions. The particle number density in the DS will thus not be conserved, and a more relevant quantity is the energy density flowing into the DS. Assuming that the DS evolves relativistically, the energy density $\rho_{\text{CFT}}$ of the DS CFT is controlled by the Boltzmann equation:
$\dot{\rho}_{\text{CFT}}=-4H(T)\rho_{\text{CFT}}+\gamma_{1}(T),$ (13)
with the energy injection rate $\gamma_{1}$ parametrized in the standard way as
$\gamma_{1}(T)=\kappa_{1}T^{5}\lambda^{2}\left(\frac{T}{\Lambda_{\text{cw}}}\right)^{2d}.$ (14)
We assume that the initial energy density of DS and the energy loss from the thermal bath are both negligible. We can find the energy density to be
$\rho(T)=\frac{3\sqrt{10}}{\sqrt{g_{*}}\pi}\frac{\kappa_{1}\lambda^{2}T_{R}^{2d}}{(2d-1)\Lambda_{\text{cw}}^{2d}}\frac{M_{\text{pl}}}{T_{R}}T^{4}.$ (15)
Here $g_{*}$ is the effective degrees of freedom of the relativistic species, and since the energy density of the DS remains small compared to the energy density of the SM, we will take the value for the SM. Notice that we have assumed that this DS production is UV-dominated, which corresponds to $d>1/2$ interactions.
By considering the $2\rightarrow\text{CFT}$ scattering process $|D_{\mu}\text{H}|^{2}\rightarrow\text{CFT}$, [24] has computed the energy injection rate coefficient $\kappa_{1}$ to be
$\kappa_{1}=\frac{a_{\mathcal{O}_{i}}a_{\mathcal{O}_{f}}}{16\pi^{5}}\frac{\Gamma(d_{i}+d_{f}-1)\Gamma(d_{i}+d_{f}-3)}{\Gamma(d_{i})\Gamma(d_{i}-1)\Gamma(d_{f})\Gamma(d_{f}-1)},$ (16)
where $a_{\mathcal{O}}$ is the normalization factor of the two-point functions, $i$ stands for the initial state and $f$ denotes the final state. For the scattering process $|D_{\mu}\text{H}|^{2}\rightarrow\text{CFT}$, we take the normalization factors to be $a_{|D_{\mu}\text{H}|^{2}}=48$ and $a_{\text{CFT}}=2\pi^{2}$. Upon collecting all the factors, the coefficient is $\kappa_{1}=0.44$ for $|D_{\mu}\text{H}|^{2}\rightarrow\mathcal{O}_{\psi}$. The energy density from the UV-dominant process will be
$\rho_{\text{CFT, UV}}=1.2\times 10^{17}\lambda^{2}T_{R}^{2.6}T^{4}\text{GeV}/\Lambda_{\text{cw}}^{3.6}.$ (17)
The IR-dominant interaction (12) will give the Higgs decay contribution [6, 7, 25]
$\gamma^{\prime}_{1}(T)$ $=n_{h}\langle\Gamma(h\rightarrow\text{CFT})E\rangle$
$=f_{\Delta}\frac{\lambda^{2}m_{h}^{2\Delta+2}v^{2}T}{\Lambda_{\text{cw}}^{2\Delta}}K_{2}\left(\frac{m_{h}}{T}\right),$ (18)
where $f_{\Delta}\equiv 4^{-\Delta}\pi^{1/2-2\Delta}\Gamma(\Delta+1/2)/(\Gamma(\Delta-1)\Gamma(2\Delta))$ and $K_{\nu}(x)$ is the modified Bessel functions of the second kind. The process roughly starts around the EW scale $\sim v$ and ends when the SM temperature reaches the Higgs mass $m_{h}$ [25]. After the temperature reaches the Higgs mass, the Higgs decouples from the thermal bath and the decay will be exponentially suppressed. We can take the limit $m_{h}\rightarrow 0$ in the thermal average computation and simplify the decay contribution:
$n_{h}\langle\Gamma(h\rightarrow\text{CFT})E\rangle=2f_{\Delta}\frac{\lambda^{2}m_{h}^{2\Delta}}{\Lambda_{\text{cw}}^{2\Delta}}v^{2}T^{3}.$ (19)
Solving the Boltzmann equation (13) and taking the temperature to be $m_{h}$ will give us
$\rho_{\text{CFT, IR}}(m_{h})$ $=\frac{\sqrt{5}M_{\text{pl}}f_{\Delta}\lambda^{2}m_{h}^{2\Delta+4}}{\pi^{3/2}\sqrt{g_{*}}v\Lambda_{\text{cw}}^{2\Delta}}\left(\frac{v^{3}}{m_{h}^{3}}-1\right)$
$=3.9\times 10^{20}\lambda^{2}m_{h}^{3.6}/\Lambda_{\text{cw}}^{3.6}\text{ GeV}^{4}.$ (20)
When $T\sim m_{h}$, the ratio between $\rho_{\text{CFT, UV}}$ and $\rho_{\text{CFT, IR}}$ is
$\frac{\rho_{\text{CFT, UV}}}{\rho_{\text{CFT, IR}}}=2.1\times 10^{-3}\left(\frac{T_{R}}{\text{GeV}}\right)^{2.6}\gtrsim 2.4\times 10^{5}.$ (21)
Thus the IR-dominant contribution is sub-dominant, and we will neglect it when estimating the DM masses.
Since the DS is strongly interacting, the DS will self-thermalize. Assuming that the thermalization is instantaneous and neglecting the loss of energy, the DS temperature at thermal equilibrium $T_{D}$ relates to the energy density as
$\rho=\frac{\pi^{2}}{30}g_{D}T_{D}^{4},$ (22)
where $g_{D}$ is the number of degrees of freedom in the DS before the phase transition. We will see at the end of this section that this is a self-consistent assumption. The UV effective degrees of freedom of (1) is $100$. Flowing to the near-IRFP will require corrections to the effective degrees of freedom, which is beyond the scope of this work. We will still take $g_{D}\sim O(10^{2})$. Combining (17) and (22), we can define the ratio between the DS temperature and the SM temperature to be
$\xi\equiv\frac{T_{D}}{T}=2.5\times 10^{4}\times\left(\frac{\lambda^{2}T_{R}^{2.6}\text{GeV}}{g_{D}\Lambda_{\text{cw}}^{3.6}}\right)^{1/4}.$ (23)
From here the DS starts to have a different temperature, but we assume the energy density to remain small so it will not affect the IR processes. We assume that the DM particles are non-relativistic and decouple from the SM, while all injected energy becomes DM masses. The DM energy density today is $\rho_{\text{DM},0}=\rho_{\text{CFT}}(T_{c})s_{0}/s_{\text{SM}}(T_{c}/\xi)$, where the SM entropy density is $s=(2\pi^{2}/45)g_{*}T^{3}$. The DM relic abundance is defined as
$\frac{\Omega_{\text{DM}}h^{2}}{0.12}=\frac{\rho_{\text{DM},0}}{\rho_{\text{crit},0}}\frac{h^{2}}{0.12}=\frac{\rho_{\text{CFT}}(T_{c})}{0.36\text{ eV}\times s_{\text{SM}}(T_{c}/\xi)}$
$=$ $3.5\times 10^{36}\frac{T_{c}}{\text{GeV}}\left(\frac{\lambda^{2}T_{R}^{2.6}\text{GeV}}{\Lambda_{\text{cw}}^{3.6}}\right)^{7/4}.$ (24)
Assuming the confining temperature to be $T_{c}\sim m_{\text{PS}}$, the DM particle mass will be
$\frac{\Omega_{\text{DM}}h^{2}}{0.12}\approx 1$
$\Rightarrow$ $m_{\text{PS}}\sim 2.9\times 10^{-37}\left(\frac{\Lambda_{\text{cw}}^{3.6}}{\lambda^{2}T_{R}^{2.6}\text{GeV}}\right)^{7/4}\text{ GeV}.$ (25)
We can see that the mass of DM particles increases with larger $\Lambda_{\text{cw}}$ and smaller $T_{R}$. By estimating the DM lifetime $\tau_{\text{DM}}\sim M_{\text{pl}}^{4}/m_{\text{PS}}^{5}>10^{28}\text{ s}$, we can setup the upper bound for the DM mass once we consider the decay into gravitons and $\gamma$-rays [24, 29]. A lower bound can be established by considering standard DM self-interaction, taking $\langle\sigma v\rangle\sim\pi/m_{\text{PS}}^{2}$. The region $m_{\text{PS}}\lesssim 10^{-1}\text{ GeV}$ will be excluded from the Bullet Cluster observations [24]. Assuming that the reheating temperature cannot exceed the inflationary Hubble constant and taking the upper and lower bounds of DM mass, we can plot the region of permitted masses, as shown in Figure 4.
Figure 4: The relation between the DM particle mass and the Wilson coefficient $\lambda$. The yellow region is the permitted mass range with respect to $T_{R}$. We take $\Lambda_{\text{cw}}$ to be the lowest value, in order to find the widest mass range. The blue regions are excluded by the Bullet Cluster observations, inflationary Hubble constant and DM lifetime.
Now we look back and argue for the self-consistency of DS self-thermalization. From the plot we can see that $\lambda\gtrsim 2\times 10^{-4}$, for the lowest value of $\Lambda_{\text{cw}}$. We assume that the self-thermalization of DS is dominated by $2\rightarrow 2$ processes. Roughly we can take the interaction rate $\Gamma\sim n_{D}\langle\sigma v\rangle$. The number density of DS is $n_{D}=\zeta(3)g_{D}T_{D}^{3}/\pi^{2}\approx 12T_{D}^{3}$. Taking the cross section $\langle\sigma v\rangle$ to be $\langle\sigma v\rangle\sim\alpha^{2}/T_{D}^{2}$, where $\alpha$ is an effective coupling, the interaction rate becomes $\Gamma=12\alpha^{2}T_{D}$. We have
$\frac{\Gamma}{H}\sim 9.1\times 10^{5}\frac{\alpha^{2}M_{\text{pl}}}{\sqrt{g_{*}}T}\left(\frac{\lambda^{2}T_{R}^{2.6}\text{GeV}}{g_{D}\Lambda_{\text{cw}}^{3.6}}\right)^{1/4}$ (26)
$\sim 9.1\times 10^{5}\frac{\alpha^{2}M_{\text{pl}}}{\sqrt{g_{*}}g_{D}^{1/4}T}\left(3.5\times 10^{36}\frac{m_{\text{PS}}}{\text{GeV}}\right)^{-1/7}$ (27)
Since we assume the self-thermalization to happen instantaneously, we can take $T$ of SM to be $T_{R}$. The ratio $\Gamma/H$ will become
$\frac{\Gamma}{H}=\frac{4.0\times 10^{17}\alpha^{2}\text{ GeV}}{\left(m_{\text{PS}}/\text{GeV}\right)^{1/7}T_{R}}.$ (28)
Thus, for a natural effective coupling $\alpha\sim O(1)$, the ratio $\Gamma/H\gtrsim O(10^{2})$, so the self-thermalization is sufficient. From Figure 4, we can also find that the IR-dominant process is not only sub-dominant but also kinetically forbidden.
@@
Discussion.—A simple gauge theory at the lower edge of the conformal window in the DS opens up new possibilities for beyond-Standard-Model (BSM) physics and DM candidates. The Dark Walker model has the potential to produce both the primordial non-Gaussianity, which is large enough to have the possibility of being detectable, and the observed relic abundance of dark matter through freeze-in. As expected, several parameters must still be fixed by hand, though their values remain within natural ranges. The non-Gaussianity signal is sensitive to the inflationary Hubble scale, the initial gauge coupling—which determines the starting energy scale of the merged fixed point—and the precise anomalous dimensions of scalar primary operators in the conformal phase. The dark matter abundance is sensitive to the critical temperature of confinement, the reheating temperature and also the energy scale of unparticle formation. These parameters may be constrained more accurately in the future through improved observations of cosmological parameters and lattice QCD studies of the spectrum in the $\text{SU}(3)$ theory with eight flavors. With such inputs, more precise predictions of the primordial non-Gaussianity and relic abundance can be achieved. This framework offers a concrete and coherent testing ground for the role of strongly coupled sectors in cosmology.
@@ Acknowledgements
I express my thanks to Riccardo Barbieri, who encouraged me very helpfully while this work was in progress. Thanks to Marco Costa, Raffaele-Tito D’Agnolo, Janet Hung, Guilherme Leite Pimentel, Alessandro Strumia and the Reviewers for useful discussions and comments.
@@
Appendix A Calculation of non-Gaussianity
In the main text, we used the EFT language of inflation to express the non-Gaussianity $f_{\text{NL}}$ as
$f_{\text{NL}}\equiv\frac{5}{18}\frac{(k_{1}k_{2}k_{3})^{2}}{(2\pi^{2})^{2}\Delta_{\zeta}^{4}(k)}\langle\zeta_{\vec{k}_{1}}\zeta_{\vec{k}_{2}}\zeta_{\vec{k}_{3}}\rangle^{\prime}\big|_{k_{1}=k_{2}=k_{3}=k}$
$=\frac{5}{18}\frac{k^{6}}{(2\pi^{2})^{2}\Delta_{\zeta}^{4}(k)}(-H_{\text{inf}})^{3}\langle\pi_{\vec{k}_{1}}\pi_{\vec{k}_{2}}\pi_{\vec{k}_{3}}\rangle^{\prime}\big|_{k_{i}=k}.$ (29)
When the scalar perturbation $\pi$ is canonically normalized as
$\pi\equiv\frac{\pi_{c}}{\sqrt{2}M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}}=\frac{\pi_{c}}{\dot{\phi}},$ (30)
the $\pi_{c}$ will behave exactly the same as $\delta\phi$. Combining with the results in [5], $f_{\text{NL}}$ can be further written as
$f_{\text{NL}}=$ $\frac{5}{18}\frac{2^{2}k^{6}}{(2\pi^{2})^{2}\Delta_{\zeta}^{4}(k)}\left(-\frac{H_{\text{inf}}}{\dot{\phi}}\right)^{3}\left(\frac{\mu}{H_{\text{inf}}}\right)^{4-2\Delta}$
$\times$ $\frac{H_{\text{inf}}^{3}}{(2k^{3})^{3}}\mathcal{B}_{\phi}\left(k,k,k\right).$ (31)
The bispectrum $\mathcal{B}_{\phi}\left(k_{1},k_{2},k_{3}\right)$ in our model is
$\mathcal{B}_{\phi}\left(k_{1},k_{2},k_{3}\right)\equiv\frac{\epsilon}{2}k_{3}^{3}U_{12}b(u)+\text{permutations},$
$U_{12}(\cdot)\equiv\frac{1}{2}\left(1-\frac{k_{1}k_{2}}{k_{12}}\partial_{k_{12}}\right)\left(\frac{1-u^{2}}{u^{2}}\partial_{u}(u\cdot)\right),$
$b(u)=\frac{2u}{u+1}\frac{1}{2-\Delta}{}_{2}F_{1}\left(1,2-\Delta;3-\Delta;\frac{2u}{u+1}\right)$
$-\left(\frac{2u}{u+1}\right)^{2(\Delta-1)}\frac{1}{\Delta-1}{}_{2}F_{1}\left(1,\Delta-1;\Delta;\frac{2u}{u+1}\right)$
$+\frac{2}{\Delta-1}\frac{2u}{u+1}{}_{2}F_{1}\left(1,1;\Delta;\frac{1-u}{1+u}\right),$ (32)
where $_{2}F_{1}$ is the hypergeometric function.